# Modeling the Deterioration of Pavement Skid Resistance and Surface Texture After Preventive Maintenance

### v1

@@ Conclusions
The results demonstrated that ensemble models such as XGBoost and Random Forest outperformed conventional methods like linear regression in predicting both skid number and macrotexture. More importantly, the proposed Transformer-based sequence-to-one modeling framework further improved predictive accuracy. Specifically, it achieved the highest $R^{2}score=0.980$ for skid number prediction and also ranked second for macrotexture prediction, which validates the effectiveness of modeling surface deterioration as a dynamic sequence prediction task. These findings support the integration of temporal learning techniques into pavement performance modeling, which enables more proactive and data-driven maintenance strategies.

### v2

@@ Conclusions
The results demonstrated that ensemble models such as XGBoost and Random Forest outperformed conventional methods like linear regression in predicting both skid number and macrotexture. More importantly, the proposed time series transformer-based modeling framework further improved predictive accuracy. Specifically, it achieved the highest $R^{2}score=0.980$ for skid number prediction and also ranked second for macrotexture prediction, which validates the effectiveness of modeling surface deterioration as a dynamic sequence prediction task. These findings support the integration of temporal learning techniques into pavement performance modeling, which enables more proactive and data-driven maintenance strategies.
Limitations
@@
In addition, several important factors that influence pavement surface deterioration, including traffic volume, the percentage of heavy vehicles, seasonal variation, and construction timing, were not included in the dataset due to limited data availability. The absence of these variables may introduce bias into the predictive model, especially in relation to traffic loading and environmental conditions. Future studies should consider incorporating these factors to improve model robustness and generalizability.
Recommendations for Future Work